The sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers.
Answer by lwsshak3(11628)

The sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers.
**
let x=larger number
let y=smaller number
..
x+y=2x-4
-x+y=-4
x-3y=-20
add
-2y=-24
y=12
x=y+4=16
ans:
larger number=16
smaller number=12